Suppose that $2000 is loaned at a rate of 7.5% , compounded semiannually. Assuming that no payments are made, find the amount owed after 7 years.

Do not round any intermediate computations, and round your answer to the nearest cent.

Answer by lwsshak3(11628)

Suppose that $2000 is loaned at a rate of 7.5% , compounded semiannually. Assuming that no payments are made, find the amount owed after 7 years.
This problem is solved by using the compound interest formula,
A=P(1+interest/period)^nbr of periods.
A=amount accummulated
P=amount loaned=$2000
Interest per period=.075/2=.0375
Number of periods = 7*2=14
A=2000(1+.0375)^14
=2000(1.0375)^14
=$3348.60
ans: The amount owed after 7 years = $3348.60